OpenStudy (anonymous):

Solve for x... √x-1=2-√x+3

OpenStudy (anonymous):

i got x=-1. Is this right?

jhonyy9 (jhonyy9):

how you have calculated please ?

jhonyy9 (jhonyy9):

can you rewrite it here now ?

OpenStudy (anonymous):

x-1=x+7-4√x+3 -8=-4√x+3..... then square both sides 64=16(x+3) 64=16x+48 -16=16x x=-1

OpenStudy (anonymous):

is this right?

jhonyy9 (jhonyy9):

how you have got in the left side -16 ,this not is right

jhonyy9 (jhonyy9):

64-48=16x

OpenStudy (anonymous):

oo sorry so it would just be 1 not negative

jhonyy9 (jhonyy9):

so than will get 16/16 = +1

jhonyy9 (jhonyy9):

yes +1 will be right sure good luck bye

OpenStudy (anonymous):

Thanks